A metal ion M forms a stable complex with the ligand X:

M + (aq) + 4 X (aq) ⇌ M (X) 4+ (aq) Kf = 1.0 ∙ 10^20

In a solution where [M +] = 0.100 M and [X] = 0.800 M before reaction, what is [X] at equilibrium?
Answer #1

Answer

[X] = 0.400M

Explanation

First consider complete formation of M(X)4+

M+(aq) + 4X(aq) ------> M(X)4+( aq) .

stoichiometrically, 1mole of M+ reacts with 4moles of X to give 1mole of M(X)4+

0.100M of M+ reacts 0.400M to give 0.100M M(X)4+

After completion of reaction

[X] = 0.800M - 0.400M = 0.400M

[M+] = 0

[M(X)4+] = 0.100M

Now, consider the dissociation equilibrium of M(X)4+

M(X)4+(aq) <-------> M+(aq) + 4X(aq)

Kd = [M+][X]4/[M(X)4+]

Kd = 1/Kf = 1/1.0 ×1020 = 1.0 ×10-20

Initial concentration

[M(X)4+] = 0.100

[M+] = 0

[X] = 0.400

chamge in concentration

[M(X)4+] = - x

[M+] = +x

[X] = +4x

Equilibrium concentration

[M(X)4+] = 0.100 - x

[ M+] = x

[X] = 0.400 + 4x

so,

x(0.400 +4x)4 / 0.100 - x = 1.0 ×10-20

We can assume 0.400 + 4x = 0.400 and 0.100 - x= 0.100 because x is small value

x(0.400)4/0.100 = 1.0 ×10-20

x 0.256 = 1.0 × 10-20

x = 3.91 × 10-20

Therefore,

[X] = 0.400 + 4( 3.91 ×10-20) = 0.400M